# A Unique Bosonic Symmetry in a 4D Field-Theoretic System

### v1

@@
Abstract: For the combined field-theoretic system of the four (3 + 1)-dimensional (4D) Abelian 3-form and 1-form gauge theories, we show the existence of a unique bosonic symmetry transformation that is constructed from the four infinitesimal, continuous and off-shell nilpotent symmetry transformations which exist for the Becchi-Rouet-Stora-Tyutin (BRST) quantized versions of the coupled (but equivalent) Lagrangian densities (that describe the present 4D field-theoretic system). The above off-shell nilpotent symmetry transformations are nothing but the BRST, co-BRST, anti-BRST and anti-co-BRST, under which, the Lagrangian densities transform to the total spacetime derivatives. The proof of the uniqueness of the above bosonic symmetry transformation operator crucially depends on the validity of all the four Curci-Ferrari (CF) type restrictions that exist on our theory. We highlight the importance of these CF-type restrictions, at various levels of our theoretical discussions, in the context of the unique bosonic transformation. We compare this observation against the backdrop of the three CF-type restrictions that appear in the requirements of the absolute anticommutativity of the specific set of nilpotent symmetries.
PACS numbers: 11.15.-q, 12.20.-m, 03.70.+k
@@ 1 Introduction
It is an undeniable truth that the ideas of global and local symmetries (see, e.g. [1-3] and references therein) have played a key role in the development of theoretical physics (right from the very inception of physics as a branch of modern-day science). In particular, the 4D standard model of elementary particle physics, where there is a stunning degree of agreement between theory and experiment, owes a great deal to the interacting non-Abelian 1-from gauge theory which is based on the local Yang-Mills gauge symmetries. One of the key theoretical methods for the covariant canonical quantization of such kinds of theories (based on the gauge symmetries) is the Becchi-Rouet-Stora-Tyutin (BRST) formalism [5-7] where the unitarity and the quantum gauge (i.e. BRST) invariance are respected together at any arbitrary order of perturbative computations (see, e.g. [8] for details).
@@
The central topic of our present endeavor is the four (3 + 1)-dimensional (4D) BRST-quantized version of the combined field-theoretic system of the free Abelian 3-form and 1-form gauge theories where we focus on the continuous as well as the discrete symmetry transformation operators and their algebraic structures. In particular, we concentrate on (i) the four continuous off-shell nilpotent symmetry transformation operators, (ii) a couple of discrete duality symmetry transformations, (iii) a unique continuous bosonic symmetry transformation operator (that is constructed from the appropriate anticommutator(s) between the above nilpotent symmetry transformation operators), and (iv) the derivation of the precise algebraic structures that are obeyed by the above continuous and discrete symmetry transformation operators. We demonstrate that there is an uncanny resemblance between the algebraic structures that are obeyed by (i) the symmetry transformation operators of our 4D BRST-quantized theory, and (ii) the de Rham cohomological operators of differential geometry (see, e.g. [9-11]). We also establish a two-to-one mapping between the symmetry transformation and cohomological operators***A set of three operators ($d,\;\delta,\;\Delta$), defined on a compact manifold without a boundary, is christened as the de Rham cohomological operators of differential geometry where $(\delta)d$ (with $\delta=\pm\,*\;d\;*,\;d=\partial_{\mu}\;(dx^{\mu}),\;\delta^{2}=0,\;d^{2}=0$) are the (co-)exterior derivatives and $\Delta=\{d,\;\delta\}$ stands for the Lalacian operator. Here the symbol * corresponds to the Hodge duality operator. These operators obey an algebra: $d^{2}=0,\;\delta^{2}=0,\;\Delta=\{d,\;\delta\},\;[\Delta,\;d]=0,\;[\Delta,\;\delta]=0$ which is popularly known as the Hodge algebra [9-11]. at the algebraic level.
@@
Our present investigation is essential and important on the following counts. First of all, in our earlier works [12-14], we have been able to show the existence of a set of four infinitesimal, continuous and off-shell nilpotent (anti-)BRST and (anti-)co-BRST symmetry transformation operators for our present 4D BRST-quantized field-theoretic combined system of the free Abelian 3-form and 1-from gauge theories. We are curious to know the algebraic structure that is obeyed by these fermionic symmetry operators in our present endeavor. Second, we have also been able to demonstrate that our 4D theory respects a couple of discrete duality symmetry transformation operators. Thus, we are very eager to know their roles in the whole gamut of the algebraic structure that is respected by various kinds of symmetry transformation operators of our theory. Third, we lay emphasis on the roles of the CF-type restrictions on our 4D BRST-quantized theory (as far as the whole beauty of the algebraic structure is concerned). Fourth, out of the six anticommutators that exist amongst the four nilpotent symmetry transformation operators, it is very interesting to show that (i) the two of them absolutely anticmmute with each-other (modulo the U(1) vector gauge symmetry-type transformations), (ii) the other two anticommutators are found to be zero provided we use the specific set of three (out of the total four) CF-type restrictions that exist on our 4D theory, and (iii) the rest of the two anticomutators define the bosonic symmetry transformation operators. Finally, we establish that, out of the above two bosonic symmetry transformation operators, only one of them is independent on the space of quantum fields where all the four CF-type restrictions are satisfied.
@@
The key theoretical contents of our present endeavor are organized as follows. In the next section, we recapitulate the bare essentials of the four nilpotent transformations that exist for our 4D BRST-quantized field-theoretic system [13,14]. In our section three, we discuss the algebraic structure of the discrete and continuous symmetry transformation operators and define a set of two bosonic symmetry transformation operators from the four nilpotent symmetry transformation operators. In particular, we point out the appearance of a specific set of three (out of the total four) CF-type restrictions in the context of the requirement of the absolute anticommutativity between (i) the (anti-)BRST symmetry transformation operators, and (ii) the (anti-)co-BRST symmetry transformation operators. Our section four deals with the bosonic symmetry transformations on (i) the basic fields of our theory, and (ii) the coupled Lagrangian densities, where we show (i) the appearance of the full set of the four CF-type restrictions, and (ii) the proof of the uniqueness of the bosonic symmetry transformation operator. Finally, in our section five, we summarize our key observations and point out the future perspective of our present investigation.
@@ 2 Preliminaries: Off-Shell Nilpotent Symmetries
[…]
@@
- $\big(\partial_{\mu}\bar{\beta}_{\nu}-\partial_{\mu}\bar{\beta}_{\nu}\big)\big(\partial^{\mu}\beta^{\nu}\big)-\,\partial_{\mu}\bar{C}_{2}\,\partial^{\mu}C_{2}\Big]-\partial_{\mu}\bar{C}\,\partial^{\mu}C,$
where the Abelian 1-form ($A^{(1)}=A_{\mu}\,dx^{\mu}$) and 3-form $A^{(3)}=\frac{1}{3!}\,A_{\mu\nu\sigma}\,(d\,x^{\mu}\wedge d\,x^{\nu}\wedge d\,x^{\sigma})$] define the free basic gauge fieldsOur 4D flat spacetime manifold is endowed with the metric tensor $\eta_{\mu\nu}$ = diag (+1, -1, -1, -1) so that the dot product between two non-null vectors $P_{\mu}$ and $Q_{\mu}$ is: $(P\cdot Q)=\eta_{\mu\nu}P^{\mu}Q^{\nu}=P_{0}Q_{0}-P_{i}Q_{i}$ where the Greek indices $\mu,\nu,\sigma...=0,1,2,3$ and the Latin indeces i,j,k=1,2,3i,j,k...=1,2,3. We choose the 4D Levi-Civita tensor $\varepsilon_{\mu\nu\sigma\rho}$ such that $\varepsilon_{0123}=+1=-\varepsilon^{0123}$ and, when two of them are contracted together, the standard relationships: $\varepsilon_{\mu\nu\sigma\rho}\varepsilon^{\mu\nu\sigma\rho}=-\,4!,\;\varepsilon_{\mu\nu\sigma\rho}\varepsilon^{\mu\nu\sigma\eta}=-\,3!\,\delta^{\eta}_{\rho},\;\varepsilon_{\mu\nu\sigma\rho}\varepsilon^{\mu\nu\eta\kappa}=-\,2!(\delta^{\eta}_{\sigma}\delta^{\kappa}_{\rho}-\delta^{\eta}_{\rho}\delta^{\kappa}_{\sigma})$, etc., are satisfied. $A_{\mu}$ and $A_{\mu\nu\sigma}$, respectively. In the Lagrangian density ${\cal L}_{(NG)}$, the set of bosonic Nakanishi-Lautrup type auxiliary fields $B,B_{1},B_{2},B_{3},B_{\mu\nu},{\cal B}_{\mu\nu}$ have been invoked for the linearization purposes. For instance, the auxiliary fields $B_{1},\,{\cal B}_{\mu\nu}$ have been utilized in the linearizations of the kinetic terms for the Abelian 3-form and 1-form gauge fields, respectively. On the other hand, the set of auxiliary fields $B,(B_{3})B_{2},B_{\mu\nu}$ have been used to linearize the gauge-fixing terms for the Abelian gauge field $A_{\mu}$, the (axial-)vector fields $(\widetilde{\phi}_{\mu})\phi_{\mu}$ (which appear in the theory because of the reducibility properties) and the Abelian 3-form gauge field $A_{\mu\nu\sigma}$, respectively. In the Faddeev-Popov (FP) ghost part of the Lagrangian density ${\cal L}_{(FP)}$, we have the basic (anti-)ghost fields: $(\bar{C}_{\mu\nu})C_{\mu\nu},\,(\bar{\beta}_{\mu})\beta_{\mu},\,(\bar{C}_{2})C_{2},(\bar{C}_{1})C_{1},\,(\bar{C})C$ with the ghost numbers: (1)+1,(2)+2,(3)+3,(1)+1,(1)+1(-1)+1,(-2)+2,(-3)+3,$(-1)+1,$ respectively. The ferminonic (anti-)ghost fields $(\bar{C})C$ are associated with the Abelian 1-form gauge field and the rest of the above basic (anti-)ghost fields are connected to the Abelian 3-from gauge field $A_{\mu\nu\sigma}$. In the above, all the basic (anti-)ghost fields are fermionic in nature except the (anti-)ghost vector fields $(\bar{\beta}_{\mu})\beta_{\mu}$ which are bosonic (i.e. $\beta_{\mu}^{2}\neq 0,\;\bar{\beta}_{\mu}^{2}\neq 0,\,\beta_{\mu}\bar{\beta}_{\nu}=\bar{\beta}_{\nu}\beta_{\mu}$, etc.) in nature. We also have fermionic auxiliary (anti-)ghost fields $(\bar{F}_{\mu})f_{\mu}$ with the ghost numbers (-1)+1 and the bosonic set of auxiliary (anti-)ghost fields $(B_{5})B_{4}$ with the ghost numbers (-2)+2, respectively. It can be checked that the following infinitesimal, continuous and off-shell nilpotent (i.e. $s_{(d)b}^{2}=0$) (co-)BRST [i.e. (dual-)BRST] symmetry transformations $s_{(d)b}$, namely;
@@
$s_{d}C_{\mu\nu}=+\,{\cal B}_{\mu\nu},\qquad s_{d}C=+\,B_{1},\quad s_{d}C_{2}=B_{4},\qquad s_{d}\bar{C}_{1}=B_{5},\qquad s_{d}F_{\mu}=+\,\partial_{\mu}B_{3},$
@@
$+$ $B_{2}\,f^{\mu}+B_{4}\,\bar{F}^{\mu}-(\partial^{\mu}\,\beta^{\nu}-\partial^{\nu}\,\beta^{\mu})\,\bar{F}_{\nu}\Big]-\partial_{\mu}\Big[B\,\partial^{\mu}C\Big].$
Corresponding to the Lagrangian density ${\cal L}_{(B)}={\cal L}_{(NG)}+{\cal L}_{(FP)}$ [cf. Eqs. (1),(2)], we have a coupled Lagrangian density ${\cal L}_{(\bar{B})}={\cal L}_{(ng)}+{\cal L}_{(fp)}$ whose non-ghost sector [i.e. ${\cal L}_{(ng)}$] and FP-ghost sector [i.e. ${\cal L}_{(fp)}$] parts (of the Lagrangian density) are
${\cal L}_{(ng)}=\dfrac{1}{2}\,B^{2}-B\,(\partial\cdot A)+\dfrac{1}{2}\,B_{2}\,\big(\partial\cdot\phi\big)-\dfrac{1}{4}\,B_{2}^{2}-\dfrac{1}{2}\,B_{3}\,\big(\partial\cdot\widetilde{\phi}\big)+\dfrac{1}{4}\,B_{3}^{2}$ (7)
@@
- $\big(\partial_{\mu}\bar{\beta}_{\nu}-\partial_{\mu}\bar{\beta}_{\nu}\big)\big(\partial^{\mu}\beta^{\nu}\big)-\,\partial_{\mu}\bar{C}_{2}\,\partial^{\mu}C_{2}\Big]-\partial_{\mu}\bar{C}\,\partial^{\mu}C,$
@@
$s_{ad}\bar{C}_{\mu\nu}=+\,\bar{\cal B}_{\mu\nu},\qquad s_{ad}\bar{C}=-\,B_{1},\qquad s_{ad}\bar{C}_{2}=-\,B_{5},\qquad s_{ad}\bar{C}_{1}=-\,B_{3},$
@@
- $B_{5}\,F^{\mu}+B_{2}\,\bar{f}^{\mu}-(\partial^{\mu}\,\bar{\beta}^{\nu}-\partial^{\nu}\,\bar{\beta}^{\mu})\,F_{\nu}\Big]-\partial_{\mu}\Big[B\,\partial^{\mu}\bar{C}\Big].$
@@ 3 Algebraic Structures: Discrete Duality and Continuous Nilpotent Symmetry Operators
In our present section, we focus on (i) the the discrete duality symmetry transformation operator, (ii) the continuous and nilpotent symmetry transformation operators, and (iii) the algebraic structures that are satisfied by them. In particular, we attach a great of importance to the emergence of the (anti-)BRST as well as the (anti-)co-BRST invariant CF-type restrictions in our present discussions. In this context, first of all, we note that the following discrete duality symmetry transformations for the basic and auxiliary fields§§§It is worthwhile to mention that our current discrete duality symmetry transformations (13) are slightly different from such transformations quoted in our earlier work [12] and these are the correct set because they satisfy accurately the algebraic structure that have been mentioned in (14) and (17).
$A_{\mu}\longrightarrow\,\mp\,\dfrac{1}{3!}\,\varepsilon_{\mu\nu\sigma\rho}\,A^{\nu\sigma\rho},\quad A_{\mu\nu\sigma}\longrightarrow\,\pm\,\varepsilon_{\mu\nu\sigma\rho}\,A^{\rho},\quad B_{\mu\nu}\to\pm\,{\cal B}_{\mu\nu},\quad{\cal B}_{\mu\nu}\to\mp\,B_{\mu\nu},$
$B\to\pm\,B_{1},\quad B_{1}\to\mp\,B,\quad B_{2}\to\pm\,B_{3},\quad B_{3}\to\mp\,B_{2},\;\;\phi_{\mu}\to\pm\,\widetilde{\phi}_{\mu},\;\;\widetilde{\phi}_{\mu}\to\mp\,\phi_{\mu},$
$C_{\mu\nu}\longrightarrow\,\pm\,\bar{C}_{\mu\nu},\quad\bar{C}_{\mu\nu}\longrightarrow\,\mp\,C_{\mu\nu},\quad\beta_{\mu}\to\mp\,\bar{\beta}_{\mu},\quad\bar{\beta}_{\mu}\to\mp\,\beta_{\mu},\quad f_{\mu}\to\pm\,\bar{F}_{\mu},$
$\bar{F}_{\mu}\to\mp\,f_{\mu},\qquad B_{4}\to\pm\,B_{5},\qquad B_{5}\to\pm\,B_{4},\qquad C\to\mp\,\bar{C},\qquad\bar{C}\to\pm\,C,$
$C_{2}\to\pm\,\bar{C}_{2},\qquad\bar{C}_{2}\to\mp\,C_{2},\qquad C_{1}\to\pm\,\bar{C}_{1},\qquad\bar{C}_{1}\to\mp\,C_{1},$ (13)
leave the Lagrangian density ${\cal L}_{(B)}={\cal L}_{(NG)}+{\cal L}_{(FP)}$ [cf. Eqs. (1),(2)] invariant. It is worthwhile to mention, at this stage, that the interplay between the continuous symmetry operators [cf. Eq. (3),(4)] and discrete symmetry operator [cf. Eq. (13)], provide the physical realization of the well-known operator relationship: $\delta=\pm\,*\,d\,*$ between the (co-)exterior derivatives $(\delta)d$ of differential geometry because we observe the following
@@
It is worthwhile to point out that the nilpotent (i.e. $\delta^{2}=0,\,d^{2}=0$) (co-)exterior derivative operators $(\delta)d$ are (i) physically realized [cf. Eq. (14)] in terms of the nilpotent (i.e. $s_{d}^{2}=0,\,s_{b}^{2}=0$), infinitesimal and continuous (dual-)BRST symmetry transformation operators $s_{(d)b}$, and (ii) the discrete duality symmetry transformation operators (13) provide the physical realizations of the Hodge duality * operator of differential geometry.

### v2

@@
Abstract: For the combined field-theoretic system of the four (3 + 1)-dimensional (4D) Abelian 3-form and 1-form gauge theories, we show the existence of a unique bosonic symmetry transformation that is constructed from the four infinitesimal, continuous and off-shell nilpotent symmetry transformations which exist for the Becchi-Rouet-Stora-Tyutin (BRST) quantized versions of the coupled (but equivalent) Lagrangian densities that describe the present 4D field-theoretic system. The above off-shell nilpotent symmetry transformations are nothing but the BRST, co-BRST, anti-BRST and anti-co-BRST, under which, the Lagrangian densities transform to the total spacetime derivatives. The proof of the uniqueness of the above bosonic symmetry transformation operator crucially depends on the validity of all the four Curci-Ferrari (CF) type restrictions that exist on our theory. We highlight the importance of these CF-type restrictions, at various levels of our theoretical discussions, in the context of the unique bosonic transformation. We compare this observation against the backdrop of the three CF-type restrictions that appear in the requirements of the absolute anticommutativity of the specific set of nilpotent symmetries.
PACS numbers: 11.15.-q; 12.20.-m; 03.70.+k
@@ 1 Introduction
It is an undeniable truth that the ideas of global and local symmetries (see, e.g. [1-3] and references therein) have played a key role in the development of theoretical physics (since the very inception of the subject of physics as a branch of modern-day science). In particular, the 4D standard model of elementary particle physics, where there is a stunning degree of agreement between theory and experiment, owes a great deal to the interacting non-Abelian 1-from (i.e. $p=1$) gauge theory which is based on the local Yang-Mills gauge symmetries. One of the key theoretical methods for the covariant canonical quantization of such kinds of theories (based on the gauge symmetries) is the Becchi-Rouet-Stora-Tyutin (BRST) formalism [5-7] where the unitarity and the quantum gauge (i.e. BRST) invariance are respected together at any arbitrary order of perturbative computations (see, e.g. [8]).
@@
The central theme of our present endeavor is the four (3 + 1)-dimensional (4D) BRST-quantized version of the combined field-theoretic system of the free Abelian 3-form***The study of the higher $p$-form ($p=2,3$...) gauge theories has become quite important because such kinds of basic fields appear in the quantum excitations of the (super)string theories which are the forefront areas of research activities in theoretical high energy physics (see, e.g. [9-11] and references therein). and 1-form gauge theories where we focus on the continuous as well as the discrete symmetry transformation operators and their algebraic structure(s). In particular, we concentrate on (i) the four continuous off-shell nilpotent symmetry transformation operators, (ii) a couple of discrete duality symmetry transformations, (iii) a unique continuous bosonic symmetry transformation operator (that is constructed from the appropriate anticommutator(s) between the above nilpotent symmetry transformation operators), and (iv) the derivation of the precise algebraic structures that are obeyed by the above continuous and discrete duality symmetry transformation operators. We demonstrate that there is an uncanny resemblance between the algebraic structures that are obeyed by (i) the discrete and continuous symmetry transformation operators of our 4D BRST-quantized theory, and (ii) the de Rham cohomological operatorsA set of three operators ($d,\;\delta,\;\Delta$), defined on a compact manifold without a boundary, is christened as the de Rham cohomological operators of differential geometry where $(\delta)d$ (with $\delta=\pm\,*\;d\;*,\;d=\partial_{\mu}\;(dx^{\mu}),\;\delta^{2}=0,\;d^{2}=0$) are the (co-)exterior derivatives and $\Delta=\{d,\;\delta\}$ stands for the Lalacian operator. Here the symbol * corresponds to the Hodge duality operator. These operators obey an algebra: $d^{2}=0,\;\delta^{2}=0,\;\Delta=\{d,\,\delta\},\;[\Delta,\,d]=0,\;[\Delta,\,\delta]=0$ which is popularly known as the Hodge algebra [12-14]. of differential geometry (see, e.g. [12-14]).
@@
Our present investigation is essential and important on the following counts. First of all, in our earlier works [15-17], we have been able to show the existence of a set of four infinitesimal, continuous and off-shell nilpotent (anti-)BRST and (anti-)co-BRST symmetry transformation operators for our present 4D BRST-quantized field-theoretic combined system of the free Abelian 3-form and 1-from gauge theories. We are curious to know the algebraic structure that is obeyed by these fermionic symmetry operators in our present endeavor. Second, we have also been able to demonstrate that our 4D theory respects a couple of discrete duality symmetry transformation operators. Thus, we are very eager to know their roles in the whole gamut of the algebraic structures that are respected by various kinds of symmetry transformation operators of our theory. Third, we lay emphasis on the roles of the CF-type restrictions on our 4D BRST-quantized theory (as far as the whole beauty of the algebraic structures are concerned). Fourth, out of the six anticommutators that exist amongst the four nilpotent symmetry transformation operators, it is very interesting to show that (i) the two of them absolutely anticommute with each-other (modulo the U(1) vector gauge symmetry-type transformationsWe observe that these specific anticommutators generate the infinitesimal and continuous symmetry transformations that change the ghost numbers of the specific set of fields (e.g. $\phi_{\mu},\;\widetilde{\phi}_{\mu},\;\beta_{\mu},\;\bar{\beta}_{\mu}$) of our 4D BRST-quantized field-theoretic system (cf. Appendix A for details).), (ii) the other two anticommutators are found to be zero provided we use the specific set of three (out of the total four) CF-type restrictions that exist on our 4D theory, and (iii) the rest of the two anticomutators define the bosonic symmetry transformation operators. Finally, we establish that, out of the above two bosonic symmetry operators, only one of them is independent on the subspace of quantum fields where all the four CF-type restrictions are satisfied.
@@
The key theoretical contents of our present endeavor are organized as follows. In the next section, we recapitulate the bare essentials of the four nilpotent transformations that exist for our 4D BRST-quantized field-theoretic system [16,17]. In our Sec. 3, we discuss the algebraic structures of the discrete and continuous symmetry transformation operators and define a set of two bosonic symmetry transformation operators from the four nilpotent symmetry transformation operators. In particular, we point out the appearance of a specific set of three (out of the total four) CF-type restrictions in the context of the requirement of the absolute anticommutativity between (i) the (anti-)BRST symmetry transformation operators, and (ii) the (anti-)co-BRST symmetry transformation operators. Our Sec. 4 deals with the bosonic symmetry transformations on (i) the basic fields of our theory, and (ii) the coupled Lagrangian densities, where we show (i) the appearance of the full set of the four CF-type restrictions, and (ii) the proof of the uniqueness of the bosonic symmetry transformation operator. Finally, in our Sec. 5, we summarize our key observations and point out the future perspective of our present investigation.
In our Appendix A, we elaborate on the specific anticommutators [cf. Eqs. (20),(21) below] and show that they are not the true bosonic symmetry transformations of our theory.
@@ 2 Preliminaries: Off-Shell Nilpotent Symmetries
We begin with the following (co-)BRST invariant Lagrangian density ${\cal L}_{(B)}={\cal L}_{(NG)}+{\cal L}_{(FP)}$ that contains the non-ghost part ${\cal L}_{(NG)}$ and the FP-ghost part ${\cal L}_{(FP)}$ as [15-17]
@@
- $\big(\partial_{\mu}\bar{\beta}_{\nu}-\partial_{\mu}\bar{\beta}_{\nu}\big)\big(\partial^{\mu}\beta^{\nu}\big)-\,\partial_{\mu}\bar{C}_{2}\,\partial^{\mu}C_{2}\Big]+\partial_{\mu}\bar{C}\,\partial^{\mu}C,$
[…]
@@
$s_{d}C_{\mu\nu}=+\,{\cal B}_{\mu\nu},\qquad s_{d}C=-\,B_{1},\quad s_{d}C_{2}=B_{4},\qquad s_{d}\bar{C}_{1}=B_{5},\qquad s_{d}F_{\mu}=+\,\partial_{\mu}B_{3},$
@@
$+$ $B_{2}\,f^{\mu}+B_{4}\,\bar{F}^{\mu}-(\partial^{\mu}\,\beta^{\nu}-\partial^{\nu}\,\beta^{\mu})\,\bar{F}_{\nu}\Big]+\partial_{\mu}\Big[B\,\partial^{\mu}C\Big].$
Corresponding to the Lagrangian density ${\cal L}_{(B)}={\cal L}_{(NG)}+{\cal L}_{(FP)}$ [cf. Eqs. (1),(2)], we have a coupled Lagrangian density ${\cal L}_{(\bar{B})}={\cal L}_{(ng)}+{\cal L}_{(fp)}$ whose non-ghost sector [i.e. ${\cal L}_{(ng)}$] and FP-ghost sector [i.e. ${\cal L}_{(fp)}$] (of the Lagrangian density ${\cal L}_{(\bar{B})}$) are
${\cal L}_{(ng)}=B\,(\partial\cdot A)-\dfrac{1}{2}\,B^{2}+\dfrac{1}{2}\,B_{2}\,\big(\partial\cdot\phi\big)-\dfrac{1}{4}\,B_{2}^{2}-\dfrac{1}{2}\,B_{3}\,\big(\partial\cdot\widetilde{\phi}\big)+\dfrac{1}{4}\,B_{3}^{2}$ (7)
@@
- $\big(\partial_{\mu}\bar{\beta}_{\nu}-\partial_{\mu}\bar{\beta}_{\nu}\big)\big(\partial^{\mu}\beta^{\nu}\big)-\,\partial_{\mu}\bar{C}_{2}\,\partial^{\mu}C_{2}\Big]+\partial_{\mu}\bar{C}\,\partial^{\mu}C,$
@@
$s_{ad}\bar{C}_{\mu\nu}=+\,\bar{\cal B}_{\mu\nu},\qquad s_{ad}\bar{C}=+\,B_{1},\qquad s_{ad}\bar{C}_{2}=-\,B_{5},\qquad s_{ad}\bar{C}_{1}=-\,B_{3},$
@@
- $B_{5}\,F^{\mu}+B_{2}\,\bar{f}^{\mu}-(\partial^{\mu}\,\bar{\beta}^{\nu}-\partial^{\nu}\,\bar{\beta}^{\mu})\,F_{\nu}\Big]+\partial_{\mu}\Big[B\,\partial^{\mu}\bar{C}\Big].$
@@ 3 Algebraic Structures: Discrete Duality and Continuous Nilpotent Symmetry Operators
In our present section, we focus on (i) the the discrete duality symmetry transformation operators, (ii) the continuous and nilpotent symmetry transformation operators, and (iii) the algebraic structures that are satisfied by them. In particular, we attach a great of importance to the emergence of the (anti-)BRST as well as the (anti-)co-BRST invariant CF-type restrictions in our present discussion. In this context, first of all, we note that the following discrete duality symmetry transformations for the basic and auxiliary fieldsIt is worthwhile to mention that our current discrete duality transformations (13) are slightly different from such transformations quoted in our earlier work [15] and these are the correct transformations because they satisfy accurately the algebraic structures that have been mentioned in equations (14) and (17).
$A_{\mu}\longrightarrow\,\pm\,\dfrac{i}{3!}\,\varepsilon_{\mu\nu\sigma\rho}\,A^{\nu\sigma\rho},\quad A_{\mu\nu\sigma}\longrightarrow\,\pm\,i\,\varepsilon_{\mu\nu\sigma\rho}\,A^{\rho},\quad B_{\mu\nu}\to\pm\,i\,{\cal B}_{\mu\nu},\quad{\cal B}_{\mu\nu}\to\pm\,i\,\,B_{\mu\nu},$
$B\to\mp\,i\,B_{1},\;\;B_{1}\to\mp\,i\,B,\;\;B_{2}\to\pm\,i\,\,B_{3},\;\;B_{3}\to\pm\,i\,B_{2},\;\;\phi_{\mu}\to\pm\,i\,\widetilde{\phi}_{\mu},\;\;\widetilde{\phi}_{\mu}\to\pm\,i\,\phi_{\mu},$
$C_{\mu\nu}\longrightarrow\,\pm\,i\,\bar{C}_{\mu\nu},\quad\bar{C}_{\mu\nu}\longrightarrow\,\pm\,i\,C_{\mu\nu},\quad\beta_{\mu}\to\pm\,i\,\bar{\beta}_{\mu},\quad\bar{\beta}_{\mu}\to\mp\,i\,\beta_{\mu},\quad f_{\mu}\to\pm\,i\,\bar{F}_{\mu},$
$\bar{F}_{\mu}\to\pm\,i\,f_{\mu},\qquad B_{4}\to\mp\,i\,B_{5},\qquad B_{5}\to\pm\,i\,B_{4},\qquad C\to\pm\,i\,\bar{C},\qquad\bar{C}\to\pm\,i\,C,$
$C_{2}\to\pm\,i\,\bar{C}_{2},\qquad\bar{C}_{2}\to\pm\,i\,C_{2},\qquad C_{1}\to\pm\,i\,\bar{C}_{1},\qquad\bar{C}_{1}\to\pm\,i\,C_{1},$ (13)
leave the Lagrangian density ${\cal L}_{(B)}={\cal L}_{(NG)}+{\cal L}_{(FP)}$ [cf. Eqs. (1),(2)] invariant. It is worthwhile to mention, at this stage, that the interplay between the continuous symmetry operators [cf. Eq. (3),(4)] and the discrete symmetry operator [cf. Eq. (13)], provide the physical realization of the well-known operator relationship: $\delta=\pm\,*\,d\,*$ between the (co-)exterior derivatives $(\delta)d$ of differential geometry because we observe the following
@@
It is worthwhile to point out that the nilpotent (i.e. $\delta^{2}=0,\,d^{2}=0$) (co-)exterior derivative operators $(\delta)d$ are (i) physically realized [cf. Eq. (14)] in terms of the nilpotent (i.e. $s_{d}^{2}=0,\,s_{b}^{2}=0$), infinitesimal and continuous (dual-)BRST symmetry transformation operators $s_{(d)b}$, and (ii) the discrete duality symmetry transformations (13) provide the physical realizations of the Hodge duality * operator of differential geometry.